OpenStudy (anonymous):

A racing dog is initially running at 13.0 , but is slowing down.How fast is the dog moving when its kinetic energy has been reduced by half?

OpenStudy (anonymous):

Formula for kinetic energy: \[ E_k = \frac{1}{2} m v^2 \] when the kinetic energy is reduced by half, we get \[ \frac{1}{2}m {v_2}^2 = \frac{1}{2}\frac{1}{2} m {v_1}^2 \] and the answer is \[ v_2 = \frac{v_1}{\sqrt{2}} \]

OpenStudy (anonymous):

ok i got like this 1/2m (13)sq, when reduced by half i solved for velocity, square root of 84.5 =9.19